Pure copper was mixed with a 23% copper alloy to produce an alloy that was 32% copper. how much of each were used to produce 44kg of 32% alloy?
Answer by checkley71(8403)

(44-X)+.23X=.32*44
44-X+.23X=14.08
-77X=14.08-44
-.77X=-29.92
X=-29.92/-.77
X=38.86
44-38.86=5.14 OZ OF PURE COPPER IS USED.
(44-38.86)+.23*38.86=14.08
5.14+8.94=14.08
14.08=14.08